# A unique $Q$-point and infinitely many near-coherence classes of ultrafilters

### v2

@@ 2 Result
Since $\langle s\cup t^{\prime}\cup\{l^{\dagger}\},x_{2}\setminus(l^{\dagger})^{+}\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}}\undertilde{C}\cap(l_{0},m^{\dagger})=D^{\dagger}$, it follows that $C_{t^{\prime}\cup\{l^{*}\}}\cap(m^{\dagger})=D^{\dagger}$ and hence $D^{\dagger}=D^{*}\cap(l_{0},m^{\dagger})$. However, both $D^{\dagger}$ and $D^{*}$ have size $\gamma$ and thus $D^{*}\subseteq(l_{0},m^{\dagger})$, which is a contradiction to the fact that $a\in D^{*}$ and $a>m^{\dagger}$. $\blacksquare$   $\dashv$
@@
Say that a $Q$-point element selects from an interval partition if it intersects each interval in exactly one point. Let $v_{0}\in E$ and $u_{0}\in\mathcal{U}$ be such that they select from $I$. Let $f$ be an order-preserving bijection from $v_{0}$ to $u_{0}$, extended to a permutation of $\omega$. Thus, for each $i\in\omega$, $f$ sends the element selected by $v_{0}$ in $[k_{i},k_{i+1})$ to the element selected by $u_{0}$ in $[k_{i},k_{i+1})$. Since $E$ and $\mathcal{U}$ are non-isomorphic, there exist $v_{1}\in[v_{0}]^{\omega}\cap E$ and $u_{1}\in[u_{0}]^{\omega}\cap\mathcal{U}$ such that $u_{1}\cap f[v_{1}]=\emptyset$. Hence, for all $i\in\omega\setminus\{0\}$:
@@
We can now finish the proof of the proposition and hence of the main theorem: Let $v\in E$, $u\in\mathcal{U}$ be given by the previous lemma for the interval partition $\{[k_{i},k_{i+1}):i\in\omega\}\cup\{[0,k_{0})\}$ constructed in the proof of Lemma 2.7. Let $G\ast H$ be any $\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}$-generic filter containing $\langle\langle s,x_{2}\rangle,\undertilde{q}^{\prime}\rangle$. By Lemma 2.7, we have that in $\mathbf{V}[G\ast H]$, whenever $\text{range}(\undertilde{a}[G\ast H])\setminus(\max s)^{+}$ intersects one of the intervals $[k_{i},k_{i+1})$, then the Mathias real $\eta$ intersects $[k_{i},k_{i+1})$ or one of the adjacent intervals $[k_{i-1},k_{i})$ or $[k_{i+1},k_{i+2})$. Since $\text{range}(\eta)$ is almost contained in $u$, the same is true for $u$ in place of $\eta$ above some $n\geq(\max s)^{+}$. Hence, $\text{range}(\undertilde{a}[G\ast H])\setminus n$ is disjoint from $v$.   $\dashv$   $\dashv$

### v3

@@ 2 Result
Since $\langle s\cup t^{\prime}\cup\{l^{\dagger}\},x_{2}\setminus(l^{\dagger})^{+}\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}}\undertilde{C}\cap(l_{0},m^{\dagger})=D^{\dagger}$, it follows that $C_{t^{\prime}\cup\{l^{*}\}}\cap(m^{\dagger})=D^{\dagger}$ and hence $D^{\dagger}=D^{*}\cap(l_{0},m^{\dagger})$. However, both $D^{\dagger}$ and $D^{*}$ have size $\gamma$ and thus $D^{*}\subseteq(l_{0},m^{\dagger})$, which is a contradiction to the fact that $a\in D^{*}$ and $a>m^{\dagger}$. This completes the proof of the Claim and Lemma 2.7 as well. $\dashv$
@@
Say that a $Q$-point element selects from an interval partition if it intersects each interval in exactly one point. Let $v_{0}\in E$ and $u_{0}\in\mathcal{U}$ be such that they select from $I$. Let $f$ be an order-preserving bijection from $v_{0}$ to $u_{0}$, extended to a permutation of $\omega$. Thus, for each $i\in\omega$, $f$ sends the element selected by $v_{0}$ in $k_{i},k_{i+1})$ to the element selected by $u_{0}$ in $k_{i},k_{i+1})$. Since $E$ and $\mathcal{U}$ are non-isomorphic, there exist $v_{1}\in[v_{0}]^{\omega}\cap E$ and $u_{1}\in[u_{0}]^{\omega}\cap\mathcal{U}$ such that $u_{1}\cap f[v_{1}]=\emptyset$. Hence, for all $i\in\omega\setminus\{0\}$:
@@
We can now finish the proof of the proposition and hence of the main theorem: Let $v\in E$, $u\in\mathcal{U}$ be given by the previous lemma for the interval partition $\{[k_{i},k_{i+1}):i\in\omega\}\cup\{[0,k_{0})\}$ constructed in the proof of Lemma 2.7. Let $G\ast H$ be any $\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}-$generic filter containing $\langle\langle s,x_{2}\rangle,\undertilde{q}^{\prime}\rangle$. By Lemma 2.7, we have that in $\mathbf{V}[G\ast H]$, whenever $\text{range}(\undertilde{a}[G\ast H])\setminus(\max s)^{+}$ intersects one of the intervals $k_{i},k_{i+1})$, then the Mathias real $\eta$ intersects $k_{i},k_{i+1})$ or one of the adjacent intervals $k_{i-1},k_{i})$ or $k_{i+1},k_{i+2})$. Since $\text{range}(\eta)$ is almost contained in $u$, the same is true for $u$ in place of $\eta$ above some $n\geq(\max s)^{+}$. Hence, $\text{range}(\undertilde{a}[G\ast H])\setminus n$ is disjoint from $v$. $\dashv$